Reports 1-1 of 1 Clear search Modify search
Detector Operation (Operations Report)
chiummo, DeRossi, Gosselin, Melo, Spinicelli - 16:53 Monday 21 September 2020 (49457) Print this report
Comment to Operator Report - Initial Status (49451)

During the operations to recover the INJ system, we had a sudden drop of power of the slave laser. This was due to a drop in the currents of both pumping diodes (see first plot). We managed to restore the usual current values by acting on the knobs in the EEroom. This was enough to recover the lock at the normal value for the slave laser and, as a consequence, for the downstream devices.

By a quick look, we did not find a possible origin in environmental probes (second plot), besides a small glitch on the ENV_CB_ELECTRIC. Since it happened to both the diodes and rather quickly, it does not seem related to the ongoing ageing of the diode2.

To be investigated.

Works in progress for INJ.

Images attached to this comment
Search Help